OpenStudy (aryana_maria2323):

Find dy, dx if f(x) = (x + 2)7x. a. the product of the quantity 7 times the natural log of the quantity x plus 2 plus 7 times x over the quantity x plus 2 and the quantity x plus 2 raised to the 7x power b. 7 times the natural log of the quantity x plus 2 plus 7 times x divided by the quantity x plus 2 c. 7x(x + 2)(7x − 1) d. 7xln(x + 2)

OpenStudy (jdoe0001):

if you could post a screenshot of the material, that'd help :)

OpenStudy (aryana_maria2323):

What do you mean? @jdoe0001

OpenStudy (jdoe0001):

tis easier than reading such and such at the such exponent, all over that, and under and below that underneath such and such :) that way we can remove any ambiguities

OpenStudy (aryana_maria2323):

Ohh well I am taking this online one second. I will screen shot it.

OpenStudy (aryana_maria2323):

This is what the answers look like.

OpenStudy (aryana_maria2323):

@jdoe0001

OpenStudy (anonymous):

hey can i answer.

OpenStudy (kayders1997):

I'm kinda confused here so 7x(x+2) is your equation?

OpenStudy (kayders1997):

I know how to differentiate

OpenStudy (anonymous):

i think @aryana_maria2323 your equation something like this\[y=(x+2)^{7x}\]

OpenStudy (kayders1997):

Hmmm?

OpenStudy (kayders1997):

It all depends on what it is its either going to be the chain rule or product rule

OpenStudy (anonymous):

take log both the side\[\log(y)=7xlog(x+2)\]now differentiate

OpenStudy (anonymous):

\[(1/y)(dy/dx)=7x+7(x+2)\]\[dy/dx=[7x+7(x+2)]y\]now put the value of 'y'

OpenStudy (aryana_maria2323):

Sorry I was having computer problems. The equation is f(x)= (x+2)^7x

OpenStudy (aryana_maria2323):

@newtonson

OpenStudy (aryana_maria2323):

Can you help @Nnesha

OpenStudy (kayders1997):

You have to take log of both sides so the 7x can go out front

OpenStudy (anonymous):

i have already solve your problem

OpenStudy (kayders1997):

Newton im confused when you take the log wouldnt it be 1/the inside?

OpenStudy (kayders1997):

Just kidding thats the natural log but I thought when you were differetiating a lo isnt it 1/alogx

OpenStudy (kayders1997):

@newtonson

OpenStudy (aryana_maria2323):

Wait can you please draw out the steps so I can see it easier?

OpenStudy (kayders1997):

Im on my phone so idk how to draw it :(

OpenStudy (kayders1997):

Oh good it calls for natural log, much easier

OpenStudy (aryana_maria2323):

Okay, @newtonson said he solved my problem already but I don't see where he got the answer.

OpenStudy (kayders1997):

I don't understand how he got it

OpenStudy (aryana_maria2323):

Can you please work it out how you would and try to help me get the answer?

OpenStudy (kayders1997):

Yes I can go on my computer so I would draw it out for you

OpenStudy (kayders1997):

If you would like

OpenStudy (aryana_maria2323):

Yes, thank you!

OpenStudy (anonymous):

your answer something look like that\[dy/dx=[7x/(x+2)^{7x}]+[7(x+2)/(x+2)^{7x}]\]

OpenStudy (aryana_maria2323):

That is none of the answers though

OpenStudy (aryana_maria2323):

I drew out all the possible answers.

OpenStudy (anonymous):

sorry,\[y=(x+2)^{7x}\]\[\ln y=7x \ln (x+2)\]\[(1/y)(dy/dx)=[7x/(x+2)]+7\ln (x+2)\]\[dy/dx=[[7x/(x+2]+[7\ln (x+2)]](x+2)^{7x}\]

OpenStudy (anonymous):

this was right solution sorry for mistake

OpenStudy (anonymous):

your option 'a' is right one

OpenStudy (kayders1997):

im getting there

OpenStudy (anonymous):

hey @aryana_maria2323 are you there

OpenStudy (aryana_maria2323):

Yes thank you @newtonson when I did it that is what I got to. Thank you again.

OpenStudy (kayders1997):

got it?

OpenStudy (aryana_maria2323):

Can you help me with a few more questions @newtonson

OpenStudy (aryana_maria2323):

Yes. Thank you for the help @kayders1997 Can you also help me with a few more questions?

OpenStudy (anonymous):

ok, if i done, i can

OpenStudy (kayders1997):

Yes I can try, I'm sorry drawing might take 10 minuntes

OpenStudy (kayders1997):

whats your question?

OpenStudy (aryana_maria2323):

One second. I was trying to answer another problem I knew how to do.

OpenStudy (aryana_maria2323):

This is my question @kayders1997 If 3x2 + y2 = 7 then evaluate the second derivative of y with respect to x when x = 1 and y = 2. Round your answer to 2 decimal places. Use the hyphen symbol, -, for negative value

OpenStudy (aryana_maria2323):

You there @kayders1997